# Approximate solutions to the shrinking core model

### v1

@@ Abstract
The shrinking core model describes the reaction of a spherical solid particle with a surrounding fluid. In this work, we revisit the SCM by deriving it from the underlying physical processes and performing a careful non-dimensionalisation, which highlights the limitations of the commonly used pseudo-steady-state approximation, particularly in liquid-solid systems where fluid and solid densities are comparable. To address these limitations, we derive approximate analytical solutions using a perturbation method that improves upon the pseudo-steady-state model. We also obtain a small-time solution capturing early transient behavior. A semi-implicit finite difference scheme is implemented to solve the full model numerically and benchmark the analytical approximations. We demonstrate that the perturbation solution provides significantly improved accuracy over the pseudo-steady-state model, especially in diffusion-limited regimes. Finally, we propose a simple fitting procedure combining the perturbation with the early-time solutions to estimate physical parameters from experimental data at minimal computational cost.
@@ 1 Introduction
The shrinking core model (SCM) is an idealised mathematical model developed to describe the reaction between a spherical solid and a surrounding reactant fluid. The reaction occurs at a sharp interface that moves inward into the solid interior as the fluid diffuses through the growing, inert, product layer. It is a classic model that has been applied to a multitude of physical problems, including: oxidation, corrosion, mineral leaching [1], anaerobic digestion, combustion [2], dehydration [3], hydrolysis [4], catalyst regeneration [5], the capture of environmental contaminants and calcination (for example through batch adsorption [6]). The present study is motivated through analysis of hydrogen storage in metal hydrides [7] however, in the following we leave the analysis general such that it may describe liquid or gas interactions with a solid core.
@@
Mathematical modelling of the SCM is generally based on the pseudo-steady state (PSS) approximation. The physical basis for the PSS is that the conversion process is much slower than the diffusion of mass to the reaction front. Mathematically this permits the neglect of the time derivative in the concentration diffusion equation, hence the concentration satisfies a steady-state equation. However, since the boundary is actually slowly moving it is not a true steady-state, since the time-dependent boundary position is present in the solution. The PSS is a well-known approximation applied in a variety of fields, such as phase change, enzyme kinetics, lubrication theory (where the time derivative is dropped from the Navier-Stokes equations) or the related thin film theory, see [8, 9, 10] for example. The problem studied in the present paper is analagous to a spherically symmetric one-phase change (Stefan problem) with a variable phase change temperature. This has been studied extensively in the analysis of nanoscale melting and nanocrystal growth [11, 12, 13, 14, 15]. The classical SCM is often attributed to Yagi and Kunii [2] and is described in detail in [16, 17, 18].
@@
The applicability of the PSS has been questioned by a number of authors. Levenspiel [17] states that his analysis is based on the assumption that the surrounding fluid is a gas, but goes on to assert that is equally applicable to a liquid. Wen [19] clearly states that the PSS is valid for most solid-gas reaction systems but not for solid-liquid systems, unless the liquid reactant concentration is very low. Lidell [20] focusses on reminding readers of the approximations and limits of the original models, which appear to have been lost over time and in particular on the range of applicability of the PSS assumption. They summarise Wen’s result [19] stating that the non-dimensional coefficient of the time derivative in the diffusion equation must be less than 0.1 (which leads to errors of the order 10%) to avoid erroneous results (in the following we will derive a similar limitation). A key result is that they point out the different system behaviour at small times (something we will also discuss later in this paper).
@@
There are clearly situations where the non-dimensional coefficient of the time derivative in the diffusion equation is not negligible, for example at early times (hence the different behaviour mentioned in [20]) or with liquid-solid systems. In these cases some form of correction must be sought. One attempt was provided by Theofanous and Lim [21]. After changing all the variables in the problem, they integrate the heat equation twice to determine an integral equation for the concentration. This permits a solution to be obtained by repeated integration (the idea being that with each further integration the solution becomes more accurate). However, the method requires an initial solution which their approach does not provide, hence they apply a PSS approximation. To make for a simpler solution form they have imposed constant concentration at either boundary which results in a linear concentration. When this is placed into the integral formulation a cubic equation is obtained (their 2nd order solution). Replacing this back into the integral the 3rd order, obviously, is a quintic. The process could be carried further following a standard iterative procedure but at the quintic stage the algebra is already becoming very cumbersome. However, there is no proof that this iterative method converges on the correct solution. Moreover, the constant concentration boundary condition corresponds to the diffusion dominated case described in Levenspiel [17] but, the diffusion controlled PSS is a cubic equation: their starting choice corresponds to the kinetic controlled limit (we will demonstrate this more clearly in subsequent sections). The solution in [21] involving the sum of their first and second order approximations, which is a cubic, may then be viewed as a somewhat inconsistent approximation to the standard model involving both kinetic and diffusion mechanisms - their governing equations neglect the kinetic reaction but their starting solution is equivalent to the kinetic solution. Liddell [20] states that the full cubic form may be obtained through elementary calculus.
@@
A review of different solid-state kinetic models can be found in [22], where a plethora of different models are briefly mathematically motivated and classified into one of the following categories: nucleation, geometrical contraction, diffusion, and those models described by the order of the underlying chemical reaction. A critical evaluation of the SCM can be found in [23]. The basic SCM involves a single particle, extending the model to multiple particles, in [23], they investigate the role of the particle size distribution and the effects of neglecting it. The main conclusion is that if the distribution is wide, with a large variance, the kinetic equations can be affected in a non-negligible way, as supported by experimental data. The interplay between multiple evolving particles is well-documented in the case of crystal growth, through the process of Ostwald ripening. A mathematical model equivalent to the SCM, but involving multiple spherical crystals is described in [24] to show that smaller particles may grow and then decay as larger particles use up the material available in the surrounding solution.
@@
The paper is organized as follows. In Section 2, we derive the shrinking core model, nondimensionalize it, and discuss the distinct time scales involved in the reaction process. We also revisit the widely used Pseudo-Steady-State (PSS) approximation, examining its derivation and highlighting the limitations of several common simplifications. In Section 3, we develop approximate analytical solutions using the perturbation method, thereby providing corrections to the PSS forms. Additionally, we analyze the small-time limit and derive explicit expressions describing the solution behavior in this regime. In Section 4, we formulate a semi-implicit finite difference numerical scheme to solve the full model. In Section 5, we compare and discuss the various analytical and numerical solutions found. We also illustrate how the analytical solutions can be used to infer physical parameter values from experimental data. Finally, Section 6 presents our conclusions.
2 The Shrinking Core Model (SCM)
@@
The SCM is an idealised model to describe the reaction and evolution of a solid spherical particle with a surrounding fluid. In a typical scenario, the fluid reacts with the solid leading to a shrinking solid core surrounded by a growing outer porous layer. The outer layer is referred to as the $\beta$-phase and the inner solid core the $\alpha$-phase. Fluid diffusing through the $\beta$-phase, allows the reaction at the outer surface of the $\alpha$-phase to continue until the solid core is fully consumed, as illustrated in Fig. 1. The model is based on the assumption that the interface separating the $\alpha$-phase from the $\beta$-phase, where the reaction takes place, is narrow, such that it may be treated as a sharp interface, $r^{*}=s^{*}(t^{*})$.
@@
Figure 1: a) Schematic of the shrinking particle with external adsorption, internal diffusion, and reaction at the receding interface. (b) Temporal evolution of the shrinking core during the process.
@@
On the outer surface the particle is in contact with the surrounding fluid. The ideal gas law gives
@@
where $P^{*}$, $T^{*}$ represent the pressure and temperature of the surrounding fluid and $R_{g}$ is the ideal gas constant. In principle, $P^{*}$ and $T^{*}$ may vary with time, for now we will assume they are constant and hence $c_{R}^{*}$ is also constant.
The fluid is transported by diffusion from the outer boundary $r^{*}=R^{*}$ to the interface $r^{*}=s^{*+}(t^{*})$, where the + superscript indicates the position immediately adjacent to the boundary in the region occupied by $c^{*}$ (superscript - refers to the adjacent point in the $\alpha$-phase). Using Fick’s law, the mass rate reaching the interface may be written as
@@
This must match the flux caused by the conversion at the interface, which is assumed proportional to the difference in concentration across the interface
@@
When dealing with the evolution of a solid interface the fluid concentration at the boundary is unknown. The solid concentration may be expressed through the Ostwald-Freundlich relation
$c^{*}(s^{*-},t^{*})=c^{*}_{eq}\exp\left(\frac{\alpha}{s^{*}}\right)\,,$ (5)
where $\alpha$ is the capillary length. Typically the capillary length is of the order nanometres and so we may assume $s^{*}(t^{*})\gg\alpha$ and hence $c^{*}(s^{*-},t^{*})\approx c^{*}_{eq}=P_{eq}/(R_{g}T)$ is the equilibrium concentration and $P_{eq}(T)$ is the threshold pressure for the reaction to take place.
Through equating the flux expressions we can now define the fluid concentration at the interface
$c^{*}(s^{*+},t^{*})=c^{*}(s^{*-},t^{*})+\frac{D}{k}\left.\frac{\partial c^{*}}{\partial r^{*}}\right|_{r^{*}=s^{*+}}\approx c^{*}_{eq}+\frac{D}{k}\left.\frac{\partial c^{*}}{\partial r^{*}}\right|_{r^{*}=s^{*+}}\,.$ (6)
@@
$c^{*}(R^{*},t^{*})$ $=c^{*}_{R}\,,\qquad D\left.\frac{\partial c^{*}}{\partial r^{*}}\right|_{r^{*}=s^{*}}=k(c^{*}(s^{*},t^{*})-c^{*}_{eq})\,.$ (7)
@@
$\left.D\frac{\partial c^{*}}{\partial r^{*}}\right|_{r^{*}=s^{*}}$ $=k\,(c^{*}(s^{*},t^{*})-c_{eq}^{*})\,,$ (13)
@@
This system holds under the following assumptions
@@
$X^{*}(t^{*})=1-\frac{s^{*3}}{R^{*3}}\,,$ (15)
such that $X^{*}(0)=0$.
@@ 2.1 Non-dimensionalisation
$r\equiv\frac{r^{*}}{R^{*}}\,,\qquad s\equiv\frac{s^{*}}{R^{*}}\,,\qquad t\equiv\frac{t^{*}}{\tau}\,,\qquad c\equiv\frac{c^{*}-c_{eq}^{*}}{c_{R}^{*}-c_{eq}^{*}}\,,$ (16)
@@
$\frac{R^{*2}\rho_{\beta}}{\tau M_{g}(c_{R,0}^{*}-c_{eq}^{*})D}\frac{ds}{dt}$ $=-\left.\frac{\partial c}{\partial r}\right|_{r=s}\,,\qquad s(0)=1.$ (20)
The coefficients of the time derivatives suggest two possible time-scales, the standard diffusion time-scale $\tau=\tau_{D}=R^{*2}/D$ and the growth time-scale $\tau=\tau_{g}=\rho_{\beta}R^{*2}/(M_{g}(c_{R,0}^{*}-c_{eq}^{*})D)$. A third time-scale, the kinetic time-scale may be identified from the boundary condition at $r=s$: if we write $D=R^{*2}/\tau_{D}$ then the coefficient in (19) becomes $k\tau_{D}/R^{*}$. The diffusion time-scale quantifies the time taken for mass to diffuse to the interface, while the kinetic time-scale $\tau_{k}=R^{*}/k$ quantifies the rate at which it reacts there.
@@
The choice of $\tau$ may be motivated by the domain of interest - do we wish to examine, growth, diffusion or kinetic behaviour? The classic Pseudo-Steady State (PSS) solution follows from setting $\tau=\tau_{g}$, that is, we work on the time-scale of the evolution and require the coefficient of the time derivative in the diffusion equation $\mathrm{Da}=\tau_{D}/\tau_{g}\ll 1$, where $\mathrm{Da}$ is a form of Damköhler number representing the ratio of reaction to diffusion rates (or equivalently diffusion to reaction time-scales). This agrees with often quoted restrictions for the PSS approximation, see [19] for example. However, it requires a further restriction which is not discussed elsewhere - the time-scale $\tau_{g}$ is chosen to permit analysis with a moving boundary. Noting that the gradient at the interface is proportional to the concentration we may write the condition (24) as
@@
Alternatively, if we choose to focus on early time behaviour we could set $\tau=\tau_{D}$ or even $\tau\ll\tau_{D}$ such that the time derivative cannot be neglected. This demonstrates that for sufficiently early times there is no PSS approximation and the system behaviour may be markedly different to that at large times - the PSS cannot be applied for times such that $t={\cal O}\left(\tau_{D}\right)={\cal O}\left(R^{*2}/D\right)$.
@@
$\left.\frac{\partial c}{\partial r}\right|_{r=s}$ $=T_{m}\,c(s,t)\,,$ (28)
$\frac{ds}{dt}$ $=-T_{m}c(s,t)\,\qquad s(0)=1\,,$ (29)
@@
$T_{m}\equiv\frac{R^{*}k}{D}=\tau_{D}/\tau_{k}\,,\qquad\mathrm{Da}\equiv\frac{M_{g}(c_{R,0}^{*}-c_{eq}^{*})}{\rho_{\beta}}=\tau_{D}/\tau_{g}\,,$ (30)
and $T_{m}$, the ratio of diffusion to reaction time-scales, is related to the Thiele modulus $h_{T}$ (such that $T_{m}=h_{T}^{2}$). The Damköhler number $\mathrm{Da}=\tau_{D}/\tau_{g}$ is the ratio of the diffusive to the growth time-scale. The denominator represents the density of the solid $\beta$-phase, while the numerator is the density of the fluid. Wen’s [19] previously stated restrictions for the PSS to hold, based on the relative densities of the phases may be viewed as stating $\mathrm{Da}\ll 1$. In gas-solid systems $\mathrm{Da}\ll 1$. However, in liquid-solid systems it is not always the case that $\mathrm{Da}$ will be small and so the size of dimensional numbers should be verified when applying the models.

### v2

@@ Abstract
The shrinking core model (SCM) describes the reaction of a solid particle with a surrounding fluid. In this work, we revisit the SCM by deriving it from the underlying physical processes and performing a careful non-dimensionalisation, which highlights the limitations of the commonly used pseudo-steady-state approximation, particularly in liquid-solid systems where fluid and solid densities are comparable. To address these limitations, we derive approximate analytical solutions using a perturbation method that improves upon the pseudo-steady-state model. We also obtain a small-time solution capturing early transient behaviour. A semi-implicit finite difference scheme is implemented to solve the full model numerically and benchmark the analytical approximations. We demonstrate that the perturbation solution provides significantly improved accuracy over the pseudo-steady-state model, especially in diffusion-limited regimes. Finally, we propose a simple fitting procedure combining the perturbation with the early-time solutions to estimate physical parameters from experimental data at minimal computational cost.
@@ 1 Introduction
The shrinking core model (SCM) is an idealised mathematical model developed to describe the reaction between a solid particle and a surrounding reactant fluid. The reaction occurs at a sharp interface that moves inward into the solid interior as the fluid diffuses through the growing, inert, porous, product layer. It is a classic model that has been applied to a multitude of physical problems, including: oxidation [1], corrosion [2], mineral leaching [3], anaerobic digestion [4], combustion [5], dehydration [6], hydrolysis [7], catalyst regeneration [8], smelting [9], glass fiber diffusion [10], in the capture of environmental contaminants and calcination (for example through batch adsorption [11]). The present study is motivated by analysis of hydrogen storage in metal hydrides [12, 13, 14, 15, 16, 17, 18], however, in the following we leave the analysis general such that it may describe liquid or gas interactions with a solid core.
@@
Mathematical modelling of the SCM is generally based on the pseudo-steady state (PSS) approximation. The physical basis for the PSS is that the conversion process is much slower than the diffusion of mass to the reaction front. Mathematically this permits the neglect of the time derivative in the concentration diffusion equation, hence the concentration satisfies a steady-state equation. However, since the boundary is actually slowly moving, it is not a true steady-state, as the time-dependent boundary position is present in the solution. The PSS is a well-known approximation applied in a variety of fields, such as phase change, enzyme kinetics, lubrication theory (where the time derivative is dropped from the Navier-Stokes equations) or the related thin film theory, see [19, 20, 21] for example. The problem studied in the present paper is analogous to a spherically symmetric one-phase change (Stefan problem) with a variable phase change temperature. This has been studied extensively in the analysis of nanoscale melting and nanocrystal growth [22, 23, 24, 25, 26]. The classical SCM is often attributed to Yagi and Kunii [5] and is described in detail in [27, 28, 29].
@@
The applicability of the PSS has been questioned by a number of authors. Levenspiel [28] states that his analysis is based on the assumption that the surrounding fluid is a gas, but goes on to assert that it is equally applicable to a liquid. Wen [30] clearly states that the PSS is valid for most solid-gas reaction systems but not for solid-liquid systems, unless the liquid reactant concentration is very low. Lidell [31] focusses on reminding readers of the approximations and limits of the original models, which appear to have been lost over time and, in particular, on the range of applicability of the PSS assumption. They summarise Wen’s result [30] stating that the non-dimensional coefficient of the time derivative in the diffusion equation must be less than 0.1 (which leads to errors of the order 10%) to avoid erroneous results (in the following we will derive a similar limitation). A key result is that they point out the different system behaviour at small times (something we will also discuss later in this paper).
@@
There are clearly situations where the non-dimensional coefficient of the time derivative in the diffusion equation is not negligible, for example at early times (hence the different behaviour mentioned in [31]) or with liquid-solid systems. In these cases some form of correction must be sought. One attempt was provided by Theofanous and Lim [32]. After changing all the variables in the problem, they integrate the heat equation twice to determine an integral equation for the concentration. This permits a solution to be obtained by repeated integration (the idea being that with each further integration the solution becomes more accurate). However, the method requires an initial solution which their approach does not provide, hence they apply a PSS approximation. To make for a simpler solution form they have imposed constant concentration at either boundary which results in a linear concentration. When this is placed into the integral formulation a cubic equation is obtained (their 2nd order solution). Replacing this back into the integral the 3rd order, obviously, is a quintic. The process could be carried further following a standard iterative procedure but at the quintic stage the algebra is already becoming very cumbersome. However, there is no proof that this iterative method converges on the correct solution. Moreover, the constant concentration boundary condition corresponds to the diffusion dominated case described in Levenspiel [28] but the diffusion-controlled PSS is a cubic equation: their starting choice corresponds to the kinetic-controlled limit (we will demonstrate this more clearly in subsequent sections). The solution in [32] involving the sum of their first and second order approximations, which is a cubic, may then be viewed as a somewhat inconsistent approximation to the standard model involving both kinetic and diffusion mechanisms - their governing equations neglect the kinetic reaction but their starting solution is equivalent to the kinetic solution. Liddell [31] states that the full cubic form may be obtained through elementary calculus.
@@
A review of different solid-state kinetic models can be found in [33], where a plethora of different models are briefly mathematically motivated and classified into one of the following categories: nucleation, geometrical contraction, diffusion, and those models described by the order of the underlying chemical reaction. A critical evaluation of the SCM can be found in [34]. In [34], the model is extended to multiple particles, and the authors investigate the role of the particle size distribution and the effects of neglecting it. The main conclusion is that if the distribution is wide, with a large variance, the kinetic equations can be affected in a non-negligible way, as supported by experimental data. The interplay between multiple evolving particles is well-documented in the case of crystal growth, through the process of Ostwald ripening. A mathematical model equivalent to the SCM, but involving multiple spherical crystals is described in [35] to show that smaller particles may grow and then decay as larger particles use up the material available in the surrounding solution.
@@
The paper is organized as follows. In Section 2, we derive the shrinking core model, non-dimensionalize it, and discuss the distinct time scales involved in the reaction process. We also revisit the widely used PSS approximation, examining its derivation and highlighting the limitations of several common simplifications. In Section 3, we develop approximate analytical solutions using the perturbation method, thereby providing corrections to the PSS forms. Additionally, we analyze the small-time limit and derive explicit expressions describing the solution behaviour in this regime. In Section 4, we formulate a semi-implicit finite difference numerical scheme to solve the full model. In Section 5, we compare and discuss the various analytical and numerical solutions found. We also illustrate how the analytical solutions can be used to infer physical parameter values from experimental data. Finally, Section 6 presents our conclusions.
2 The Shrinking Core Model
@@
The SCM is an idealised model to describe the reaction and evolution of a solid particle with a surrounding fluid. In a typical scenario, the fluid reacts with the solid, leading to a shrinking solid core surrounded by a growing outer porous layer. The outer layer is referred to as the $\beta$-phase and the inner solid core as the $\alpha$-phase. Fluid diffusing through the $\beta$-phase, allows the reaction at the outer surface of the $\alpha$-phase to continue until the solid core is fully consumed, as illustrated in Fig. 1. The model is based on the assumption that the interface separating the $\alpha$-phase from the $\beta$-phase, where the reaction takes place, is narrow, such that it may be treated as a sharp interface, $r^{*}=s^{*}(t^{*})$.
@@
Figure 1: (a) Schematic of the shrinking particle showing external adsorption, internal diffusion, and reaction at the receding interface. (b) Temporal evolution of the shrinking core during the reaction process.
@@
On the outer surface, the particle is in contact with the surrounding fluid. The ideal gas law gives
@@
where $P^{*}$ and $T^{*}$ represent the pressure and temperature of the surrounding fluid, and $R_{g}$ is the ideal gas constant. In principle, $P^{*}$ and $T^{*}$ may vary with time. For now, we will assume they are constant, and hence $c_{R}^{*}$ is also constant. This is a usual setup in experiments on metal hydrides (see, e.g., [18]).
The fluid is transported by diffusion from the outer boundary $r^{*}=R^{*}$ to the interface $r^{*}=s^{*+}(t^{*})$, where the $+$ superscript indicates the position immediately adjacent to the boundary in the region occupied by $c^{*}$ (the superscript -- refers to the adjacent point in the $\alpha$-phase). Using Fick’s law, the mass rate reaching the interface may be written as
@@
This must match the flux caused by the conversion at the interface, which is assumed to be proportional to the difference in concentration across the interface:
@@
When dealing with the evolution of a solid interface, the fluid concentration at the boundary is unknown. The solid concentration may be expressed through the Ostwald–Freundlich relation:
$c^{*}(s^{*-},t^{*})=c^{*}_{\rm eq}\exp\left(\frac{\alpha}{s^{*}}\right)\,,$ (5)
where $\alpha$ is the capillary length. Typically, the capillary length is of the order of nanometres, and so we may assume $s^{*}(t^{*})\gg\alpha$ and hence $c^{*}(s^{*-},t^{*})\approx c^{*}_{\rm eq}=P_{\rm eq}/(R_{g}T)$ is the equilibrium concentration, and $P_{\rm eq}(T)$ is the threshold pressure for the reaction to take place.
Through equating the flux expressions we can now define the fluid concentration at the interface,
$c^{*}(s^{*+},t^{*})=c^{*}(s^{*-},t^{*})+\frac{D}{k}\left.\frac{\partial c^{*}}{\partial r^{*}}\right|_{r^{*}=s^{*+}}\approx c^{*}_{\rm eq}+\frac{D}{k}\left.\frac{\partial c^{*}}{\partial r^{*}}\right|_{r^{*}=s^{*+}}\,.$ (6)
@@
$c^{*}(R^{*},t^{*})$ $=c^{*}_{R}\,,\qquad D\left.\frac{\partial c^{*}}{\partial r^{*}}\right|_{r^{*}=s^{*}}=k(c^{*}(s^{*},t^{*})-c^{*}_{\rm eq})\,.$ (7)
@@
$\left.D\frac{\partial c^{*}}{\partial r^{*}}\right|_{r^{*}=s^{*}}$ $=k\,(c^{*}(s^{*},t^{*})-c_{\rm eq}^{*})\,,$ (13)
@@
This system holds under the following assumptions:
@@
$X(t^{*})=1-\frac{s^{*3}}{R^{*3}}\,,$ (15)
such that $X(0)=0$.
@@ 2.1 Non-dimensionalisation
$r\equiv\frac{r^{*}}{R^{*}}\,,\qquad s\equiv\frac{s^{*}}{R^{*}}\,,\qquad t\equiv\frac{t^{*}}{\tau}\,,\qquad c\equiv\frac{c^{*}-c_{\rm eq}^{*}}{c_{R}^{*}-c_{\rm eq}^{*}}\,,$ (16)
@@
$\frac{R^{*2}\rho_{\beta}}{\tau M_{g}(c_{R,0}^{*}-c_{\rm eq}^{*})D}\frac{ds}{dt}$ $=-\left.\frac{\partial c}{\partial r}\right|_{r=s}\,,\qquad s(0)=1.$ (20)
The coefficients of the time derivatives suggest two possible time-scales, the standard diffusion time-scale $\tau=\tau_{D}=R^{*2}/D$ and the growth time-scale $\tau=\tau_{g}=\rho_{\beta}R^{*2}/(M_{g}(c_{R,0}^{*}-c_{\rm eq}^{*})D)$. A third time-scale, the kinetic time-scale may be identified from the boundary condition at $r=s$: if we write $D=R^{*2}/\tau_{D}$ then the coefficient in (19) becomes $k\tau_{D}/R^{*}$. The diffusion time-scale quantifies the time taken for mass to diffuse to the interface, while the kinetic time-scale $\tau_{k}=R^{*}/k$ quantifies the rate at which it reacts there.
@@
The choice of $\tau$ may be motivated by the domain of interest - do we wish to examine, growth, diffusion or kinetic behaviour? The classic pseudo-steady-state solution follows from setting $\tau=\tau_{g}$, that is, we work on the time-scale of the evolution and require the coefficient of the time derivative in the diffusion equation $\mathrm{Da}=\tau_{D}/\tau_{g}\ll 1$, where $\mathrm{Da}$ is a form of Damköhler number representing the ratio of reaction to diffusion rates (or equivalently diffusion to reaction time-scales). This agrees with often quoted restrictions for the PSS approximation, see [30] for example. However, it requires a further restriction which is not discussed elsewhere - the time-scale $\tau_{g}$ is chosen to permit analysis with a moving boundary. Noting that the gradient at the interface is proportional to the concentration we may write the condition (24) as
@@
Alternatively, if we choose to focus on early time behaviour we could set $\tau=\tau_{D}$ or even $\tau\ll\tau_{D}$ such that the time derivative cannot be neglected. This demonstrates that for sufficiently early times there is no PSS approximation and the system behaviour may be markedly different to that at large times - the PSS cannot be applied for times such that $t=\mathcal{O}(\tau_{D})=\mathcal{O}(R^{*2}/D)$.
@@
$\left.\frac{\partial c}{\partial r}\right|_{r=s}$ $=\mathrm{T}_{m}\,c(s,t)\,,$ (28)
$\frac{ds}{dt}$ $=-\mathrm{T}_{m}c(s,t)\,\qquad s(0)=1\,,$ (29)
@@
$\mathrm{T}_{m}\equiv\frac{R^{*}k}{D}=\tau_{D}/\tau_{k}\,,\qquad\mathrm{Da}\equiv\frac{M_{g}(c_{R,0}^{*}-c_{\rm eq}^{*})}{\rho_{\beta}}=\tau_{D}/\tau_{g}\,,$ (30)
and $\mathrm{T}_{m}$, the ratio of diffusion to reaction time-scales, is related to the Thiele modulus $h_{T}$ (such that $\mathrm{T}_{m}=h_{T}^{2}$). The Damköhler number $\mathrm{Da}=\tau_{D}/\tau_{g}$ is the ratio of the diffusive to the growth time-scale. The denominator represents the density of the solid $\beta$-phase, while the numerator is the density of the fluid. Wen’s [30] previously stated restrictions for the PSS to hold, based on the relative densities of the phases may be viewed as stating $\mathrm{Da}\ll 1$. In gas-solid systems $\mathrm{Da}\ll 1$. However, in liquid-solid systems it is not always the case that $\mathrm{Da}$ will be small and so the size of dimensional numbers should be verified when applying the models.